5: An object of mass m is projected with initial velocity v0 at an angle of θ degrees

to the horizontal. discuss the motion(position, velocity, acceleration) of the object

throughout the period of the flight. With appropriate choice of values for param￾eters (m, v0 and θ), plot the velocity and acceleration as functions of time on the

same graph. Discuss these graphs. Assume that the surrounding air offers no resistance


1
Expert's answer
2022-05-10T13:21:22-0400

"t=\frac{2v_0\sin \theta}g,"

"h=\frac{v_0^2\sin^2\theta}{2g},"

"l=\frac{v_0^2\sin2\theta}{g}."